# OFFSIDE: Benchmarking Unlearning Misinformation in Multimodal Large Language Models

### v1

@@ Abstract
Advances in Multimodal Large Language Models (MLLMs) intensify concerns about data privacy, making Machine Unlearning (MU), the selective removal of learned information, a critical necessity. However, existing MU benchmarks for MLLMs are limited by a lack of image diversity, potential inaccuracies, and insufficient evaluation scenarios, which fail to capture the complexity of real-world applications. To facilitate the development of MLLMs unlearning and alleviate the aforementioned limitations, we introduce OFFSIDE, a novel benchmark for evaluating misinformation unlearning in MLLMs based on football transfer rumors. This manually curated dataset contains 15.68K records for 80 players, providing a comprehensive framework with four test sets to assess forgetting efficacy, generalization, utility, and robustness. OFFSIDE supports advanced settings like selective unlearning and corrective relearning, and crucially, unimodal unlearning (forgetting only text data). Our extensive evaluation of multiple baselines reveals key findings: (1) Unimodal methods (erasing text-based knowledge) fail on multimodal rumors; (2) Unlearning efficacy is largely driven by catastrophic forgetting; (3) All methods struggle with "visual rumors" (rumors appear in the image); (4) The unlearned rumors can be easily recovered and (5) All methods are vulnerable to prompt attacks. These results expose significant vulnerabilities in current approaches, highlighting the need for more robust multimodal unlearning solutions. The code is available at https://github.com/zh121800/OFFSIDE.
@@ 1 Introduction
With the rapid development and widespread application of multimodal large language models (MLLMs), models pre-trained on large-scale corpora can quickly adapt to various downstream tasks, such as visual question answering (Antol et al., 2015; Goyal et al., 2017; Pang et al., 2025), visual understanding (Sugiyama et al., 2007; Guo et al., 2016; Li et al., 2024c; Xu et al., 2025a), and reasoning (Johnson et al., 2017; Perez et al., 2018; Li et al., 2025a; Wang et al., 2025). However, during both the pretraining and post-training phases, unwanted content, such as private information and rumors, may be included, which could lead to the leakage of personal privacy and the spread of misinformation. These raise concerns about the security of MLLMs (Chen et al., 2025). Machine Unlearning (MU) (Wang et al., 2024b; Deng et al., 2025) has been proposed to address these ethical and security concerns in MLLMs, aiming to eliminate the influence of unwanted data and its effects on model performance without requiring retraining from scratch, while also complying with legal frameworks (Dang, 2021).
Given that MLLMs integrate knowledge across multiple modalities, a growing line of work has begun to study MU within multimodal contexts (Liu et al., 2024c; Xu et al., 2025b; Dontsov et al., 2024; Li et al., 2024b). However, existing benchmarks commonly rely on generative models (e.g., Arc2Face (Papantoniou et al., 2024), Flux (Labs, 2024), StyleGAN2 (Karras et al., 2020), and DuelGAN (Wei et al., 2022)) to synthesize images, neglecting harmful cues embedded in the visual modality and risking the introduction of biases that diverge from real-world distributions (Westerlund, 2019; Dolhansky et al., 2020). Moreover, existing benchmarks fail to support selectively removing specific information in an image while preserving unrelated information, typically deleting all text linked to a given image (Cheng et al., 2023). In addition, they pay little attention to the downstream effects of unlearning on other post-training procedures, such as continual learning (Wang et al., 2024a), despite its importance for evaluating practical utility  (Van de Ven & Tolias, 2019). Taken together, these limitations result in an incomplete assessment of multimodal unlearning, underscoring the need for a comprehensive benchmark tailored to MLLMs.
Table 1: Benchmark Comparison. OFFSIDE is the first to support (1) multi-image entity association (group images for each player), (2) selective unlearning of private attributes while preserving shared knowledge, (3) corrective relearning, a continual learning setting, and (4) Unimodal unlearning (unlearn through only pure text data).
@@
Type Source Entity Complete Selective Corrective Unimodal
@@
MUSE (Shi et al., 2024) - - -
TOFU (Maini et al., 2024) - - -
MMUBench (Li et al., 2024b) Real World MIKE (Li et al., 2024a) multiple
MLLMU-Bench (Liu et al., 2024c) Synthetic Arc2Face (Papantoniou et al., 2024) Single
PEBench (Xu et al., 2025b) Synthetic Flux (Labs, 2024) multiple
CLEAR (Dontsov et al., 2024) Synthetic StyleGAN2 (Karras et al., 2020) multiple
@@
In this view, we propose OFFSIDE, a benchmark inspired by football transfer market rumors, aimed at simulating diverse real-world scenarios. OFFSIDE consists of 15.68K manually created Vision-Question-Answer (VQA) pairs, with 7.84K dedicated to the multimodal unlearning and 7.84K for the unimodal unlearning. It features four distinct datasets: Forget Set, Retain Set, Test set and Relearn Set, each designed to evaluate specific aspects of unlearning methods, including unlearning efficacy, generalizability, model utility, and robustness, across both multimodal and unimodal settings. A comprehensive comparison between previous benchmarks and OFFSIDE is shown in Table 1. In OFFSIDE, each player (entity) is linked to a set of images containing both private information (e.g., transfer records) and shared information (e.g., age, height, and name). The diverse text-image connections are designed for the selective unlearning setting, which only removes the private information of the target rumor and saves the shared ones. Meanwhile, we also simulate a corrective relearning framework to investigate the sources of the unlearning abilities of the tested methods.
As shown in Figure 1. We define four real-world scenarios for evaluation: the Complete Unlearning: setting tests whether unlearning methods can completely remove all knowledge related to specific entities; the Selective Unlearning setting evaluates the ability to accurately erase particular image-text associations without affecting other learned information; the Corrective relearning setting simulates a continual learning framework to examine whether previously unlearned rumors can be successfully recovered after post-training; and the Unimodal Unlearning setting assesses whether existing LLM unlearning methods can seamlessly adapt to the multimodal context of MLLMs.
@@
Five baselines are evaluated on four distinct datasets. Our comprehensive evaluation spans a variety of tasks, including classification, generation, MM-Bench (Liu et al., 2024a), and GPT evaluator. After extensive experiments, we observe several key findings, each stemming from our specially designed experimental settings, highlighting the advantages of our datasets in providing a realistic and diverse evaluation for the multimodal unlearning task. Our key contributions are as follows:
  • We propose OFFSIDE, a novel multimodal unlearning benchmark that provides four real-world scenarios (Complete Unlearning, Selective Unlearning, Corrective Relearning, and Unimodal Unlearning), demonstrating the practical value of multimodal unlearning in real-world applications.
  • OFFSIDE provides a comprehensive evaluation of unlearning, assessing forget quality, model utility, and robustness. It is the first to explore the relationship between unlearning methods and continual learning, offering valuable insights for future research.
  • After extensive experiments with five representative MU methods, we provide empirical insights into MLLM MU: (1) Unimodal unlearning methods are ineffective at unlearning multimodal rumors. (2) The unlearned rumors can be easily recovered through continual learning. (3) All baselines fail to unlearn visual rumors (rumors appear in the image). (4) All baselines are vulnerable to prompt attacks (e.g., classification task). (5) Unlearning efficacy is largely driven by catastrophic forgetting. These findings reveal key limitations of contemporary MLLM methods, thereby motivating advances specific to multimodal learning and underscoring OFFSIDE as an essential, realistic benchmark.
@@ 2 Related Work
LLM Machine Unlearning. Existing benchmarks in LLM MU have been used to test unlearning in various contexts, such as elimination of personal identification data (Patil et al., 2023), copyright protection (Eldan & Russinovich, 2023) and harmful content removal (Lu et al., 2022). Gradient Ascent (GA) (Yao et al., 2024b) was introduced to optimize the model parameters so as to maximize the removal of targeted information from the training data. However, GA often degrades performance on the retained set. Subsequent methods, including gradient descent (GD) (Liu et al., 2022), KL-based objectives (Yao et al., 2024a; Liu et al., 2024b), and “I don’t know” (IDK) losses (Maini et al., 2024), were proposed to exert finer control over the outputs of unlearned models and to mitigate collateral damage. Additionally, Negative Preference Optimization (NPO) (Zhang et al., 2024) reframes LLM unlearning as a preference-optimization problem.
MLLM Machine Unlearning.  Li et al. (2024b) introduced a token-level KL-divergence loss for MU in MLLMs, marking a pioneering attempt to apply MU in such settings. MMUNLEARNER (Huo et al., 2025) proposes a selective unlearning approach that removes visual patterns tied to a specific entity while preserving the corresponding textual knowledge within the LLM backbone. CLEAR (Dontsov et al., 2024) extends TOFU (Maini et al., 2024) by pairing personas with textual biographies and AI-generated images, and MLLMU-Bench (Liu et al., 2024c) targets the removal of private information. PULSE (Kawakami et al., 2025) extends MLLMU-Bench to include pretrained knowledge unlearning as well as continual forgetting. PEBench (Xu et al., 2025b) is the first to categorize multimodal unlearning targets into identities and events, where the targets can reside in both the textual and visual modalities. Nevertheless, the generated entities and events are overly simplistic, leading to an almost perfect unlearning effect (close to 100%), which complicates the accurate assessment of each method’s strengths and weaknesses. The benchmarks mentioned above merely evaluate the unlearned model, overlooking its potential to integrate with other post-training methods, such as continual learning. OFFSIDE is proposed to tackle the mentioned problems above: all of the images are from real-world football players, and both the image and text may contain harmful information. Additionally, we track the model’s general capabilities at different stages on MM-Bench (Liu et al., 2024a) to ensure that the unlearning process does not compromise its overall performance.
3 OFFSIDE: Unlearn football Transfer markt rumors And Relearn facts
We introduce OFFSIDE inspired by rumors in the football transfer market, where both images and accompanying text may contain inaccurate information, potentially leading the model to propagate misinformation. OFFSIDE includes 640 images of 80 football players from 20 clubs, with each image paired with 8 shared and 6 private VQA pairs. All images and their corresponding texts in our dataset are manually curated from real-world sources and cross-checked by multiple annotators, ensuring a robust benchmark for evaluating existing unlearning methods. An overview of OFFSIDE, including its data construction pipeline and evaluation procedure, is shown in Figure 2.
@@ 3.1 Models and Data Splitting
We consider a standard machine unlearning setup, with specific designs tailored for MLLMs. For all experiments, We use Qwen2.5-VL-3B and Qwen2.5-VL-7B  (Bai et al., 2025) as the base models. Let $\mathcal{D}$ denote the full dataset, which is partitioned into four disjoint subsets: $\mathcal{D}_{\textbf{forget}}$ (Forget Set), $\mathcal{D}_{\textbf{retain}}$ (Retain Set), $\mathcal{D}_{\textbf{test}}$ (Test set), and $\mathcal{D}_{\textbf{relearn}}$ (Relearn Set). In the first stage, we obtain the vanilla model by fine-tuning the pretrained MLLMs with supervision (SFT) on $\mathcal{D}_{\textbf{forget}}\cup\mathcal{D}_{\textbf{retain}}$. During the subsequent unlearning stage, various unlearning methods are applied, with access to $\mathcal{D}_{\textbf{forget}}\cup\mathcal{D}_{\textbf{retain}}$. After unlearning, we assess the model’s utility by retraining on $\mathcal{D}_{\textbf{relearn}}$ to reintroduce corrected information, simulating a continual learning setting.
The subsets can be further categorized into private and shared sets. The private set consists of QA pairs that are unique to a single image, whereas the shared set contains QA pairs that are common across multiple images of the same player. In the Selective Forget Private Information setting, we transfer the shared set of rumor-related images from $\mathcal{D}_{\textbf{forget}}$ to $\mathcal{D}_{\textbf{retain}}$, thereby simulating a more realistic scenario in which only private information is removed while shared attributes are preserved.
All four subsets are employed to enable a comprehensive evaluation. Specifically:
@@
  • $\mathcal{D}_{\textbf{relearn}}$ assesses the effectiveness of unlearning methods in conjunction with post-training procedures, highlighting the model’s ability to adapt in a continual learning context. In this study, we specifically examine whether the knowledge that was unlearned can be recovered through the relearning process.
@@
Figure 2: Overview of the OFFSIDE framework. The MLLM is first fine-tuned on the forget and retain set to obtain the vanilla model, during which it learns the rumors associated with each player. Various unlearning methods are then applied on forget set to obtain the unlearned model. After unlearning, the model is fine-tuned on the relearn set to simulate a continual learning setting. Performance is evaluated on four distinct subsets after both the unlearning and relearning stages.
3.2 Data Construction:
@@
Unlike previous MLLM machine unlearning benchmarks (Dontsov et al., 2024; Liu et al., 2024c; Xu et al., 2025b), all of the data in OFFSIDE is manually curated. The data construction process consists of three stages:
@@
Image Curation: We manually selected 80 players from 20 Premier League clubs using Google search  [^1] All images are manually selected from https://www.google.com/imghp?hl=en. For each player, we curated the following image sets: three images representing different club periods ($\mathcal{D}_{\textbf{retain}}$), one image related to a transfer rumor ($\mathcal{D}_{\textbf{forget}}$), three test images ($\mathcal{D}_{\textbf{test}}$), and one image for relearning the facts ($\mathcal{D}_{\textbf{relearn}}$). Here, $\mathcal{D}_{\textbf{test}}$ is an augmented version of $\mathcal{D}_{\textbf{retain}}$ (see Figure 2).
@@
Text Description Curation: For each image, we constructed 14 QA pairs, comprising 6 that capture private information (e.g., the player’s market value, transfer fee) and 8 that capture shared information (e.g., the player’s height, birthdate). This design is specifically tailored for a selective unlearning setting, where the aim is to forget certain rumors or private information while retaining shared facts. Additionally, we created pure-text versions of each question-answer pair to test whether existing LLM unlearning methods can be directly extended to MLLMs.
Curating Text-Image Connections: For each player, there are 8 images representing different stages of their career at various clubs. Three images are used for training, three are reserved for testing, and one image contains a rumor about the player’s transfer history, representing the forgotten data that we aim to remove. The remaining images are used for the continual learning setting, simulating the correction of rumors. Each image is associated with 14 VQA pairs, of which 8 capture shared information such as the player’s name, nationality, and height, which are consistent across all images of the player. The remaining 6 pairs correspond to private information, specifically the player’s transfer details, with each image containing unique transfer information.

### v2

@@ Abstract
Advances in Multimodal Large Language Models (MLLMs) intensify concerns about data safety, making Machine Unlearning (MU), the selective removal of harmful/private information, a critical necessity. However, existing MU benchmarks for MLLMs are limited by a lack of image diversity, coarse-grained unlearning target, and insufficient evaluation scenarios, which fail to capture the complexity of real-world applications. To facilitate the development of MLLMs unlearning and alleviate the aforementioned limitations, we introduce OFFSIDE, a novel benchmark for evaluating misinformation unlearning in MLLMs. This manually curated dataset contains 15.68K records for 80 players, providing a comprehensive framework with four test sets to assess forgetting efficacy, generalization, utility, and robustness. OFFSIDE supports advanced unlearning targets, such as fine-grained unlearning and visual rumor removal. Our extensive evaluation of multiple baselines not only extends key findings from LLM MU to MLLM MU: (1) unlearned rumors can be easily recovered through relearning and (2) all methods are vulnerable to prompt attacks, but also introduces novel insights in the context of MLLM: (1) unimodal methods fail to handle multimodal rumors, (2) unlearning efficacy is primarily driven by catastrophic forgetting statistically, and (3) all methods struggle with visual rumors (rumors embedded in images). These results expose significant vulnerabilities in current approaches, highlighting the need for more robust multimodal unlearning solutions.
OFFSIDE: Benchmarking Unlearning Misinformation in Multimodal Large Language Models
Hao Zheng1111Equal contribution.   Zirui Pang2111Equal contribution.   Ling li3   Zhijie Deng3   Yuhan Pu3 Zhaowei Zhu4  Xiaobo Xia5  Jiaheng Wei3111Equal contribution. [^2] Corresponding author. 1Harbin Institute of Technology  2University of Illinois Urbana-Champaign 3The Hong Kong University of Science and Technology (Guangzhou) 4BIAI, ZJUT & D5Data.ai  5National University of Singapore 2022211977@stu.hit.edu.cn,   jiahengwei@hkust-gz.edu.cn
@@ 1 Introduction
With the rapid development and widespread application of multimodal large language models (MLLMs), models pre-trained on large-scale corpora can quickly adapt to various downstream tasks, such as visual question answering (antol2015vqa; goyal2017making), visual understanding (sugiyama2007methods; guo2016deep; li2024georeasoner), and reasoning (johnson2017clevr; perez2018film; li2025recognition). However, during both the pretraining and post-training phases, unwanted content, such as private information and harmful rumors, may be included, which could lead to the leakage of personal privacy and the spread of misinformation. These raise concerns about the security of MLLMs (chen2025safeeraserenhancingsafetymultimodal). Machine Unlearning (MU) (wang2024llm; deng2025guard) has been proposed to address these ethical and security concerns in MLLMs, aiming to eliminate the influence of unwanted data and its effects on model performance without requiring retraining from scratch, while also complying with legal frameworks (dang2021right). Given that MLLMs integrate knowledge across multiple modalities, a growing line of work has begun to study MU within multimodal contexts (liu2024protecting; xu2025pebench; dontsov2024clear; li2024single). However, existing benchmarks commonly rely on generative models (e.g., Arc2Face (papantoniou2024arc2face)) to synthesize images, risking the introduction of biases that diverge from real-world distributions (westerlund2019emergence; dolhansky2020deepfake; pang2025label) and neglecting harmful cues embedded in the visual modality. Moreover, existing benchmarks fail to support a fine-grained unlearning target which removes specific information in an image while preserving unrelated information, typically deleting all text linked to a given image (cheng2023can). In addition, they pay little attention to the downstream effects of unlearning on other post-training procedures, such as continual learning (wang2024comprehensive). Taken together, these limitations result in an incomplete assessment of multimodal unlearning, underscoring the need for a comprehensive benchmark tailored to MLLMs.
@@
Type Source Entity Complete Fine-grained Corrective Unimodal
@@
MUSE (shi2024muse) - - -
TOFU (maini2024tofu) - - -
MMUBench (li2024single) Real World MIKE (li2024mike) multiple
MLLMU-Bench (liu2024protecting) Synthetic Arc2Face (papantoniou2024arc2face) Single
PEBench (xu2025pebench) Synthetic Flux (flux2024) multiple
CLEAR (dontsov2024clear) Synthetic StyleGAN2 (karras2020analyzing) multiple
@@
Table 1: Benchmark Comparison. OFFSIDE is the first to support (1) multi-image entity association (group images for each player), (2) fine-grained unlearning targets, (3) corrective relearning, and (4) unimodal unlearning (unlearn through only pure text data).
In this view, we propose OFFSIDE, a benchmark inspired by visual rumors, aimed at simulating diverse real-world scenarios. It features four distinct datasets: Forget Set, Retain Set, Test set and Relearn Set, each designed to evaluate specific aspects of unlearning methods, including unlearning efficacy, generalizability, model utility, and robustness, across both uni and multi-modal settings. A comprehensive comparison between previous benchmarks and OFFSIDE is shown in Table 1.
Experiments are conducted under four real-world scenarios (as shown in Figure 1): the Complete Unlearning setting, which is similar to previous benchmarks (liu2024protecting; xu2025pebench; dontsov2024clear); the Fine-grained Unlearning setting, which evaluates the ability to accurately erase particular image-text associations without affecting other benign information; the Corrective Relearning setting, which examines whether previously unlearned rumors can be successfully recovered after post-training; and the Unimodal Unlearning setting, which assesses whether unimodal unlearning methods can seamlessly adapt to the multimodal context of MLLMs.
@@
We evaluate five classic unlearning baselines across four distinct datasets. Our comprehensive evaluation spans a variety of tasks, including classification, generation, MM-Bench (liu2024mmbench), and GPT evaluator. After extensive experiments, we observe several key findings, each stemming from our specially designed experimental settings, highlighting the advantages of our datasets in providing a realistic and diverse evaluation for the multimodal unlearning task. Our key contributions are as follows:
  • We propose OFFSIDE, a novel multimodal unlearning benchmark that provides four real-world scenarios (Complete Unlearning, Fine-grained Unlearning, Corrective Relearning, and Unimodal Unlearning), demonstrating the practical value of multimodal unlearning in real-world applications.
  • OFFSIDE provides a comprehensive framework for unlearning targets and evaluation in MLLM MU, assessing forgetfulness quality, model utility, and robustness. To the best of our knowledge, we are the first to raise the problem of unlearning deceptive visual rumors and fine-grained targets.
  • After extensive experiments, we not only extend key findings from LLM to MLLM: (1) unlearned rumors can be recovered through relearning and (2) all methods are vulnerable to prompt attacks, but also introduce novel insights in the context of MLLM: (1) unimodal methods fail to address multimodal rumors, (2) unlearning efficacy is statistically driven by catastrophic forgetting, and (3) all methods struggle with visual rumors, where rumors are embedded in images. Theses findings highlight significant limitations of current MLLM methods, underscoring the need for targeted advancements in multimodal unlearning.
@@ 2 Related Work
MLLM Machine Unlearning. MMUBench li2024single is a benchmark specifically designed to facilitate the unlearning of real-world entities. It introduces a token-level KL-divergence loss for model unlearning (MU) in multimodal large language models (MLLMs), representing a pioneering effort to apply MU in this context. CLEAR (dontsov2024clear) extends TOFU (maini2024tofu) by pairing personas with textual biographies and AI-generated images, and MLLMU-Bench (liu2024protecting) targets the removal of private information. MMUNLEARNER (huo2025mmunlearner) proposes a selective unlearning approach that removes visual patterns tied to a specific entity while preserving the corresponding textual knowledge within the LLM backbone. PULSE (kawakami2025pulse) extends MLLMU-Bench to include pretrained knowledge unlearning as well as continual forgetting. PEBench (xu2025pebench) is the first to categorize multimodal unlearning targets into identities and events, where these targets can span both textual and visual modalities. However, the generated entities and events are overly simplistic, resulting in an almost perfect unlearning effect (close to 100%), which complicates the accurate evaluation of each method’s strengths and weaknesses. The aforementioned benchmarks primarily assess the unlearned model, neglecting its potential integration with other post-training methods, such as continual learning.
In contrast, OFFSIDE addresses these issues by using images of real-world football players, where both the images and texts may contain harmful information. Additionally, we monitor the model’s overall capabilities at different stages using MM-Bench (liu2024mmbench) to ensure that the unlearning process does not degrade its general performance.  [^1] Extra related work and discussion are provided in the Appendix A.
3 OFFSIDE: Unlearn Football Transfer Market Rumors and Relearn Facts
We present OFFSIDE, a benchmark inspired by visual rumors of football players, where both images and accompanying text may contain inaccuracies that could lead the model to propagate misinformation. OFFSIDE consists of 640 images representing 80 football players from 20 different clubs. Each image is paired with 8 shared and 6 private VQA pairs. A detailed overview of the OFFSIDE dataset, including its data construction pipeline and evaluation procedure, is depicted in Figure 2.
@@ 3.1 Models and Data Splitting
We consider a standard machine unlearning setup, with specific designs tailored for MLLMs. For all experiments, We use Qwen2.5-VL-3B and Qwen2.5-VL-7B  (Qwen2.5-VL) as the base models. Let $\mathcal{D}$ denote the full dataset, which is partitioned into four disjoint subsets: $\mathcal{D}_{\textbf{forget}}$ (Forget Set), $\mathcal{D}_{\textbf{retain}}$ (Retain Set), $\mathcal{D}_{\textbf{test}}$ (Test set), and $\mathcal{D}_{\textbf{relearn}}$ (Relearn Set).
In the first stage, we obtain the vanilla model by fine-tuning the pretrained MLLMs with supervision (SFT) on $\mathcal{D}_{\textbf{forget}}\cup\mathcal{D}_{\textbf{retain}}$. During the subsequent unlearning stage, various unlearning methods are applied, with access to $\mathcal{D}_{\textbf{forget}}\cup\mathcal{D}_{\textbf{retain}}$. After the unlearning process, we evaluate the model’s utility by retraining it on $\mathcal{D}_{\textbf{relearn}}$, which reintroduces the corrected information. Specifically, $\mathcal{D}_{\textbf{relearn}}$ contains the corrected versions of the same rumors, providing updated data about the same entity.
In the Fine-grained unlearning setting, the previously mentioned subsets are further categorized into private and shared sets, simulating a more realistic scenario where only private information is removed, while shared attributes are retained. The private sets consist of QA pairs that are unique to a specific image, whereas the shared sets contain QA pairs that are common across multiple images of the same player. All four subsets are employed to enable a comprehensive evaluation. Specifically:
@@
  • $\mathcal{D}_{\textbf{relearn}}$ is used to evaluate the effectiveness of unlearning methods in conjunction with other post-training procedures, specifically assessing the model’s ability to recover knowledge that was previously unlearned during the relearning process. [^2] The evaluation on $\mathcal{D}_{\textbf{relearn}}$ is similar to $\mathcal{D}_{\textbf{retain}}$. Retain set is frozen at this stage.
@@
Figure 2: Overview of the OFFSIDE framework. The MLLM is first fine-tuned on the forget and retain set to obtain the vanilla model. Various unlearning methods are then applied on forget set to obtain the unlearned model. After unlearning, the model is fine-tuned on the relearn set to correct the rumors. Performance is evaluated on four distinct subsets after both the unlearning and relearning stages.
3.2 Visual Rumors
In the context of MLLMs unlearning, there are various unlearning targets. Previous benchmarks have primarily focused on pure-text targets (liu2024protecting; dontsov2024clear; kawakami2025pulse), where private or rumor-related information is typically embedded in the text, often neglecting the visual targets embedded within images. In contrast, OFFSIDE includes confusing visual transfer information in each image of the Forget Set. This creates a more complex and realistic scenario.
3.3 Fine-grained Unlearning
In OFFSIDE, each player is linked to a set of images containing both private information (e.g., transfer records) and shared information (e.g., age, height, and name). The diverse text-image connections are designed for the fine-grained unlearning setting, which only removes the private information of the target rumor and saves the shared ones. Previous research primarily focuses on coarse-grained unlearning, treating all information related to an individual equally (e.g., forgetting all details about a player, singer, or politician). However, this approach is unrealistic, as in real-life scenarios, we don’t require a model to forget all information about a specific individual. Such a drastic unlearning would severely impair the model’s usability, as we still want the model to retain its general cognitive abilities after the unlearning process. In this view, the goal of unlearning should be to selectively forget sensitive, rumor-related, or private information about an individual while maintaining the model’s overall functionality.
3.4 Data Construction
@@
All OFFSIDE data is manually curated. The data construction process consists of two stages:
@@
Image Curation: We manually selected 80 players from 20 Premier League clubs using Google search  [^3] All images are manually selected from https://www.google.com/imghp?hl=en. For each player, we curated the following image sets: three images representing different club periods ($\mathcal{D}_{\textbf{retain}}$), one image related to a visual transfer rumor ($\mathcal{D}_{\textbf{forget}}$), three test images ($\mathcal{D}_{\textbf{test}}$), and one image for relearning the facts ($\mathcal{D}_{\textbf{relearn}}$). Here, $\mathcal{D}_{\textbf{test}}$ is an augmented version of $\mathcal{D}_{\textbf{retain}}$. $\mathcal{D}_{\textbf{relearn}}$ is a corrected version of $\mathcal{D}_{\textbf{forget}}$(new data).
@@
Text Description Curation: For each image, we constructed 14 QA pairs, comprising 6 that capture private information (e.g., the player’s market value, transfer fee) and 8 that capture shared information (e.g., the player’s height, birthdate). This design is specifically tailored for a fine-grained unlearning setting, where the aim is to forget certain rumors(private information) while retaining shared facts. Additionally, we created pure-text versions of each question-answer pair to test whether existing LLM unlearning methods can be directly extended to MLLMs.